OpenStudy (anonymous):

A certain radioactive isotope has a half-life of 15 days. If one is to make a table showing the half-life decay of a sample of this isotope from 32 grams to 1 gram; list the time (in days, starting with t = 0) in the first column and the mass remaining (in grams) in the second column, which type of sequence is used in the first column and which type of sequence is used in the second column?

imqwerty (imqwerty):

so mass will keep on decreasing and it will get 1/2 in evry 15 days so at t=0 mass=32gms after 15 days mass=32/2 =16gms after 30 days mass=16/2 = 8gms after 45 days mass=8/2=4gms after 60days mass=4/2gms=2gms after 75days mass=2/2gms =1gm

OpenStudy (anonymous):

is that all? i thought there was more o_o

OpenStudy (jack1):

to assist with the 2nd part of the question: In an Arithmetic Sequence the difference between one term and the next is a constant. A Geometric sequence is a mathematical sequence consisting of a sequence in which the next term originates by multiplying the predecessor with a constant, better known as the common ratio. An exponential function is a specialised form of geometric sequence

imqwerty (imqwerty):

oh yea the 1st column is following an arithmetic sequence 15,30,45,60,75 with common difference of 15 the 2nd column is following a geometric sequence 32,16,8,4,2,1 with common ratio 1/2